# ”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder

### v1

@@ 5.2.1. In-Situ Compulsion Prevention
For trigger avoidance type of compulsions, such as avoiding eye contact with people that induce intrusive thoughts, the system can render visual augmentations that encourages interactions with the trigger. For instance, when avoidance of triggers is detected, the system can render a gray-scale filter on everything the user sees except the triggering person to encourage eye contact, and turn off the filter when eye contact is maintained for a preset length of time as an exposure practice.
@@
5.2.2. In-Situ Cognitive Support
Besides physical methods, mental support inspired by ACT and cognitive therapy is another important strategy that effectively adjust people’s attitudes on their obsessions. Self-managment technologies should thus consider employing contextualized suggestions to facilitate acceptance of thoughts, once mental distress or compulsion is detected. For instance, the technology can guide the user to acknowledge the uncertainty to the user’s feared consequences after encountering the trigger. Combined with professional expertise, recent advancement in LLMs have enabled such context-aware intervention to support general mental well-being (Kumar et al., 2024; Nie et al., 2024). Moreover, value-based intervention has been suggested and implemented in prior work (Khan et al., 2011; Xu et al., 2022) as a ‘wake up call’ to remind the user of committing to their values. For example, when the user with contamination OCD is compulsively washing hands, a photo of their family will be displayed on their phone to encourage them to stop, if they identified ’spending time with family’ as a value. Alternatively, we can design in-situ cognitive support based on cognitive therapy. Such technology can guide the user via conversations to challenge their obsessions, such as reframing obsessive thoughts to more hopeful ones (Sharma et al., 2023), and analyzing the likelihood of feared consequences to reject obsessive thoughts (Yule and Whittal, 2017). Although the feedback design was proposed for in-situ self-management, we believe they can be effectively adopted for out-of-session practices to reinforce therapy outcome.
@@ 5.3. Addressing Privacy Needs for People with OCD
Privacy concerns have been discussed in prior literature about personal information sensing and tracking (Harari, 2020; Torous et al., 2018). To preserve users’ privacy, the tracking technology should explicitly inform them about what kind of data are being collected, what kind of data are inferred from collected data, and who has access to collected data (Harari, 2020). When it comes to AI models, on-device inference is preferred to minimize the risk of privacy invasion (Lee et al., 2019). However, due to technology limitation, many AI services require sending user data to remote servers for inference (OpenAI, 2024). In the context of OCD, researchers should pay attention to sensitive information that can exacerbates OCD users’ mental health issues. Users should be prompted to identify the type of sensitive data, such as audio or video involving fear, and taboo thoughts (e.g., sexual and violent thoughts). The system should then actively detect sensitive information in collected data and obfuscate them, such as blurring out sensitive elements in videos and photos (Zhang et al., 2024; Tseng et al., 2024), and reframing sensitive text to less disturbing alternatives (OpenAI, 2024) before sending them out for inference. The collection of biometric data (e.g., physiological data for mental state inference) can also pose privacy concerns due to potential unauthorized use of those data to infer personal identity (Marciano, 2019). Future OCD technology should adopt privacy-enhancing approaches such as feature transformation (Melzi et al., 2024; Nandakumar and Jain, 2015) to lower the risk.
@@ 5.4. Reducing Frictions in Social Interactions
Our studies suggest that OCD poses challenge to smooth social interactions due to the need for compulsions. Prior work also revealed that people with OCD experienced more interpersonal problems primarily due to certain worries and rumination than people without OCD (Solem et al., 2015). Due to the low public awareness of OCD, people with OCD usually need to endure the awkwardness caused by their symptoms during social interactions. Disclosing one’s OCD status can earn accommodation or understanding from people (P10 in Study 1). However, not all people with OCD are willing to disclose their mental health status, particularly when their OCD involves taboo thoughts (e.g., P2’s sexual OCD in Study 1). For those who are still building symptom management skills, additional support is needed to reduce frictions caused by OCD in everyday social interactions.
With current technology, researchers can explore the use of motivational interviewing via conversational agent (Miller and Rollnick, 2012; Oewel et al., 2024) to encourage the user to keep engaged with social interactions. Technology that delivers unobtrusive intervention for emotion regulation can also be used to address the mental distress in such situation, e.g., reducing anxiety level via haptic feedback simulating a low heart rate (Costa et al., 2016, 2019; Kelling et al., 2016). That said, technology can provide limited assistance to satisfy OCD people’s social needs without improved public awareness of OCD, and appropriate social support. More stakeholders in the field, including people with OCD, mental health professionals, policy makers, should be involved to enable a more inclusive and accessible environment for people with OCD.
@@ 5.5. Technology Adoption Concerns
The low adoption rate of technology by people with OCD revealed in both studies raises concerns about the practicality of our proposed technology design opportunities. Prior research in accessibility has pointed out key factors that affect the adoption of assistive technologies, including ease of use (Scherer, 2017; Phillips and Zhao, 1993), trust (Scherer, 2017), self-consciousness (Parette and Scherer, 2004; Scherer, 2017), and social acceptance (Shinohara and Wobbrock, 2011; Scherer, 2017). These factors all can potentially affect the adoption rate of suggested technology for OCD in-situ intervention. In addition, mixed-reality is still in its early stages (Cummings et al., 2023), and options for consumer level physiological sensing technology is limited. Given this situation, mainstream technology such as smartphone- and smartwatch-based interventions might remain the most practical solution due to their broad user base, high social acceptance (Shinohara and Wobbrock, 2011) and increasingly powerful sensors and processors that can support our proposed intervention designs (Dai et al., 2021; Incel and Bursa, 2023).
@@ 6. Limitations
Our research has several limitations. First, while participants in Study 1 represented diverse OCD themes, it is unclear whether our findings apply to less common OCD themes, e.g., relationship OCD. Future work should include a broader range of OCD themes to explore differences in management strategies and technology needs. Second, we recruited participants with mild to moderate OCD at the time of the study (Study 1). Although nine had experienced severe OCD before therapy, their self-management strategies and needs may vary with symptom severity. Hence, our findings should be generalized to people with severe OCD with caution.
@@ 7. Conclusion
Our research contributed the first in-depth investigation to understand OCD people’s specific experiences and their technology needs for symptom self-management. Through an interview study with 10 participants with diverse OCD symptoms, we identified their individualized triggers, compulsions, properties that affect trigger intensity, and revealed their self-management strategies. With another interview study with seven OCD therapists, we revealed the nuances in treatment strategies in complement to those used by people with OCD. Combining the two studies, we identified gaps in OCD self-management despite currently available resources, and discuss design opportunities for personalized symptom tracking, in-situ OCD intervention, and addressing privacy and social needs that are unique to OCD.
@@ Appendix A Appendix
Table A.1. Detailed personal experience of OCD (obsessions, triggers and compulsions) of reported by participants in Study 1. Only representative triggers are listed in cases where participants reported more than three triggers.
Themes Sub-themes Example codes
Trigger characterization Spectrum of triggers sound of utensil bumping, before bed time, doubt
Sensory attributes color, sticky texture, liquid more gross than solid
Affordances door knob, bar in bus, trash can with a lid
Spatial attributes proximity to trigger, amount of triggers, position/pose of trigger
Certainty of contamination origins unknown food source, who touched the food, people they trust
Other factors affecting trigger intensity mental stress, self-care, physical stress
Connecting triggers to compulsions Direct trigger interaction jiggling the door lock, moving stuff from heating vents
Consequence mitigation cleaning, changing outfit, body scanning for symptoms
Assisted compulsions seeking reassurance, asking whether telling truth
Effect of environment familiarity on compulsions more intense with closed people, compulsion less prominent in public
OCD symptom management strategies Exposure practice in daily life contaminate with battery, maintaining eye contact with triggers
Accepting & challenging obsessive thoughts uncertainty mindset, thought diffusion, focusing on values
Reducing the urge for compulsions self reminding compulsion is unnecessary, compulsion not working from experience, make compulsion more difficult
Reassurance logging taking picture, changing to 24hr time
Relaxation breathing skills, muscle relaxation
Self-compassion reflections from success, celebrate breakthrough day
External mental health support not judging shameful behaviors, family accommodating to OCD
Challenges in OCD symptom management OCD-induced social barriers OCD interrupt social interactions, downplay to get trust
Inconvenient symptom logging & tracking no self-tracking, digital journal on symptom tracking
Table A.2. Themes, Sub-themes and Example Codes generated from Study 1
Themes Sub-themes Example codes
OCD diagnosis and assessment Identify triggers interview, self-monitor triggering events
Identify compulsion untypical compulsion, mental compulsion, duration on compulsion
Psychoeducation to gain insights Psychoeducation as standalone session explain how therapy works, explain what to expect
Improve behavior awareness by homework self-monitor intrusive thoughts, self-monitor compulsion
Active treatment via ERP Building exposure hierarchy using subjective experience SUDS, ask extent to stay in anxiety, ask extent to conduct rituals
Examine trigger properties identify fear together, fluid process, identify wiggle room
Designing diverse exposures in vivo exposures, imaginal exposures, material exposures
Exposure Practices distress detection, ask mental compulsion, practice until decreased SUDS
Issues with ERP only focus on behavioral change, mental compulsion to compensate behavioral
Making long-term progress Using relaxation techniques suitably only after exposure, only for long exposure, relaxation causes hopelessness
Promoting out-of-session success via homework people not good at tracking , identify homework barriers, afraid to do homework
Technology recommendation for self-symptom management OCDefeat, low motivation, mindfulness app, concern about privacy
Table A.3. Themes, Sub-themes and Example Codes generated from Study 2

### v2

@@ 5.2.1. Augmented Reality-Based Compulsion Prevention Support
For trigger avoidance type of compulsions, such as avoiding eye contact with people that induce intrusive thoughts (P2), the system can render visual augmentations that encourages interactions with the trigger. For instance, when avoidance of triggers is detected, the system can render a gray-scale filter on everything the user sees except the triggering person to encourage eye contact, and turn off the filter when eye contact is maintained for a preset length of time as an exposure practice.
@@
5.2.2. LLM-based Cognitive Support
In addition to physical methods, mental strategies inspired by ACT and cognitive therapy are essential for helping individuals reshape their attitudes toward obsessions. LLMs can be leveraged to offer personalized suggestions based on these strategies when mental distress or compulsions are detected. For instance, the technology can guide the user to embrace uncertainty following a triggering event. Moreover, value-based intervention techniques (Khan et al., 2011; Xu et al., 2022) can serve as a ‘wake up call’ to remind the user of committing to their values—for instance, displaying a photo of a loved one on the user’s phone to interrupt compulsive hand-washing, if the user has identified “spending time with family” as a core value. Alternatively, we can design intervention based on thought reframing—guiding the user via conversations to challenge their obsessive thoughts and reassess the likelihood of feared outcomes (Sharma et al., 2023; Yule and Whittal, 2017).
However, generating effective, context-aware responses requires carefully crafted prompts, which can be difficult without domain-specific expertise (Priyadarshana et al., 2024). Although prior work has explored LLM-based support for general CBT (Kian et al., 2024; Huang et al., 2024), no existing study has focused on adapting these techniques for OCD-specific contexts. Future work should explore not only prompt engineering for pre-trained LLMs but other techniques such as retrieval-augmented generation (Lewis et al., 2020; Guo et al., 2024) to integrate OCD-specific knowledge and user history for more targeted and effective intervention.
@@ 5.3. Addressing Privacy Needs for People with OCD
Privacy concerns have been discussed in prior literature about personal information sensing and tracking (Harari, 2020; Torous et al., 2018). To preserve users’ privacy, the tracking technology should explicitly inform them about what kind of data are being collected, what kind of data are inferred from collected data, and who has access to collected data (Harari, 2020). When it comes to AI models, on-device inference is preferred to minimize the risk of privacy invasion (Lee et al., 2019). However, due to technology limitation, many AI services require sending user data to remote servers for inference (OpenAI, 2024). In the context of OCD, researchers should pay attention to sensitive information that can exacerbates OCD users’ mental health conditions. Users should be prompted to identify the type of sensitive data, such as audio or video involving fear, and taboo thoughts (e.g., sexual and violent thoughts). The system should then actively detect sensitive information in collected data and obfuscate them, such as blurring out sensitive elements in videos and photos (Zhang et al., 2024; Tseng et al., 2024), and reframing sensitive text to less disturbing alternatives (OpenAI, 2024) before sending them out for inference. The collection of biometric data (e.g., physiological data for mental state inference) can also pose privacy concerns due to potential unauthorized use of those data to infer personal identity (Marciano, 2019). Future OCD technology should adopt privacy-enhancing approaches such as feature transformation (Melzi et al., 2024; Nandakumar and Jain, 2015) to lower the risk.
@@ 5.4. Reducing Frictions in Social Interactions
Our study suggest that OCD poses challenge to smooth social interactions due to the need for compulsions. Prior work also revealed that people with OCD experienced more interpersonal problems primarily due to certain worries and rumination than people without OCD (Solem et al., 2015). Due to the low public awareness of OCD, people with OCD usually need to endure the awkwardness caused by their symptoms during social interactions. Disclosing one’s OCD status can earn accommodation or understanding from people (P10). However, not all people with OCD are willing to disclose their mental health status, particularly when their OCD involves taboo thoughts (e.g., P2’s sexual OCD). For those who are still building symptom management skills, additional support is needed to reduce frictions caused by OCD in everyday social interactions.
With current technology, researchers can explore the use of motivational interviewing via conversational agent (Miller and Rollnick, 2012; Oewel et al., 2024) to encourage the user to keep engaged with social interactions. Technology that delivers unobtrusive intervention for emotion regulation can also be used to address the mental distress in such situation (Costa et al., 2016, 2019; Kelling et al., 2016). That said, technology can provide limited assistance to satisfy OCD people’s social needs without improved public awareness of OCD, and appropriate social support. More stakeholders in the field, including people with OCD, mental health professionals, policy makers, should be involved to enable a more inclusive and accessible environment for people with OCD.
@@ 5.5. Technology Adoption Concerns
The low adoption rate of technology by people with OCD revealed in our study raises concerns about the practicality of our proposed technology design opportunities. Prior research in accessibility has pointed out key factors that affect the adoption of assistive technologies, including ease of use (Scherer, 2017; Phillips and Zhao, 1993), trust (Scherer, 2017), self-consciousness (Parette and Scherer, 2004; Scherer, 2017), and social acceptance (Shinohara and Wobbrock, 2011; Scherer, 2017). These factors all can potentially affect the adoption rate of suggested technology for OCD just-in-time intervention. In addition, mixed-reality is still in its early stages (Cummings et al., 2023), and options for consumer level physiological sensing technology is limited. Given this situation, mainstream technology such as smartphone- and smartwatch-based interventions might remain the most practical solution due to their broad user base, high social acceptance (Shinohara and Wobbrock, 2011) and increasingly powerful sensors and processors that can support our proposed intervention designs (Dai et al., 2021; Incel and Bursa, 2023).
@@ 6. Limitations
Our research has several limitations. While our OCD participants represented diverse OCD themes, it is unclear whether our findings apply to less common OCD themes, e.g., relationship OCD. Future research should include a broader range of OCD themes to explore differences in management strategies and needs. In addition, we recruited participants with mild to moderate OCD at the time of the study. Although nine had previously experienced severe OCD symptoms, their self-management strategies and needs may vary with symptom severity. Hence, our findings should be generalized to people with severe OCD with caution.
Although interviews with therapists enriched our understanding of treatment frameworks and common challenges in symptom management, their perspectives reflect clinical knowledge rather than lived experience. While prioritizing the views of people with OCD, the involvement of OCD professionals in our study may have influenced how we organized and interpreted participant responses. Future research should involve more people with OCD from diverse background as co-researchers to ensure that lived experience remains central throughout the research process and to counterbalance clinical assumptions.
We acknowledge the potential compensation issue in our study. We provided reasonable compensation to both participant groups but offered OCD professionals a higher hourly rate to facilitate recruitment given their limited availability. This imbalance in compensation may raise ethical concerns. Future research should strive to provide comparable compensation for similar tasks across participant groups, based on time commitment and participation burden, to ensure fairness in the recruitment process.
Lastly, our work is a need-finding study (Patnaik and Becker, 1999) aimed at surfacing challenges in OCD self-management and identifying technology opportunities. In the future work, we will implement and evaluate the interventions proposed in this study to validate their feasibility and effectiveness.
@@ 7. Conclusion
Our research contributed the first in-depth investigation to understand OCD people’s specific experiences and their technology needs for just-in-time symptom management. Through an interview study with 10 participants with diverse OCD symptoms and seven OCD professionals, we identified individualized triggers and compulsions, explored properties that influence trigger intensity, and discussed how these symptoms can be recognized and tracked. We also examined the self-management strategies adopted by OCD participants, highlighting both alignment with and deviations from professional-recommended practices. Finally, we identified gaps in OCD self-management despite currently available resources, and discuss design opportunities for personalized symptom tracking, just-in-time OCD intervention, and addressing privacy and social needs that are unique to OCD.
Acknowledgements.
We sincerely thank all participants who generously shared their time, experiences, and insights with us. This work would not have been possible without their willingness to contribute. We also thank Ting Zhou and Linda Zeng for their valuable input on the study design during the early stages of the project.
@@ Appendix A Appendix
Table A.1. Detailed personal experience of OCD (obsessions, triggers and compulsions) reported by participants. For participants who reported more than three triggers, only representative examples are listed.